OpenStudy (anonymous):

if root 2 sin (60-A)=1, then A is what?

OpenStudy (anonymous):

Step 1: Divide both sides by \(\sqrt{2}\)

OpenStudy (anonymous):

sin (60-A) = 1/sqrt2 60-A = 45 degrees A = 15 degrees 60-A could also be 135 degrees and greater it depends on the range of answers you require

OpenStudy (anonymous):

sqrt(2) *sin (60-A)=1 ==> sin (60-A)=1/sqrt(2) ===> 60-A=45===> A=15